# The minimal periodicity for integral bases of pure number fields

### v1

@@ 3. Main theorems
Fix $p^{e}\,\|\,n$ and $v=v_{p}$. For $f(x)=x^{n}-a$ we distinguish the two standard branches:
(i) The unit branch $v(a)=0$. Choose $u\in\mathbb{Z}_{p}^{\times}$ with $u^{n}\equiv a\pmod{p}$ and set $\varphi_{1}=x-u$. Write the $\varphi_{1}$–adic development
$f(x)=\sum_{i=0}^{n}a_{1,i}\,\varphi_{1}(x)^{i},\qquad a_{1,0}=u^{n}-a,\quad a_{1,i}=\binom{n}{i}u^{n-i}\ (i\geq 1).$
By Kummer’s theorem one has $v\big{(}\binom{n}{i}\big{)}\geq e-v(i)$ with equality for $i=p^{k}$ ($0\leq k\leq e$). Hence the lower convex hull $N^{-}_{\varphi_{1}}(f)$ contains the points $(p^{k},e-k)$ and its principal sides are exactly the segments joining $(p^{k-1},e-(k-1))$ to $(p^{k},e-k)$, each of height $1$ (Lemma A.1). [^1] The leftmost point $(0,v(u^{n}-a))$ lies on or above the first side; it never creates a principal side of height >1>1. Inductively, passing from order t1t-1 to order $t$ consumes one extra pp–adic digit of $a$ (Lemma A.2); thus all order–$t$ residual polynomials, for $1\leq t\leq e$, and hence the entire pp–type up to order $e$, are determined by $a\bmod p^{t+1}$ and in particular by $a\bmod p^{e+1}$.
(ii) The ramified branch $v(a)>0$. With $\varphi_{1}=x$ the order–$1$ polygon is the single side from $(0,v(a))$ to $(n,0)$, so $f$ is pp–regular at order $1$ and no pp–denominators occur in the shape at such $p$ (this is the convention $r_{p}(a)=-1$ and $d_{p}(a)=0$ used in Theorem 2.1). This is immediate from the binomial expansion and the definition of residual coefficients.
At higher orders, the pure binomial form forces a height–one phenomenon at each successive principal side up to order $e$: for $p^{e}\parallel n$ the broken line $N^{-}_{\varphi_{1}}(f)$ (in the $v(a)=0$ branch) runs from $(0,v(u^{n}-a))$ to $(1,e)$, then to $(p,e-1),\dots,(p^{e-1},1)$ and finally to $(p^{e},0)$, so each principal side drops by exactly one unit of height. Thus passing from order t1t-1 to order $t$ introduces exactly one further division by $p$ before reduction, and the order–$t$ residual polynomial depends only on $a\bmod p^{t+1}$. This pattern is standard in the pure case.
By induction on $t=1,\dots,e$ we conclude that the entire pp–type (hence the pp–integral basis, and so the local pp–shape) is determined by the residue class of $a$ modulo $p^{e+1}$ (see Lemma A.2). This proves (i).
We show that $p^{e}$ precision does not suffice in general. Take units at $p$:
$a=1+p^{e}u,\qquad a^{\prime}=1+p^{e+1}u^{\prime},\qquad p\nmid u,u^{\prime}.$
For odd $p$, the lifting–the–exponent formula gives $v_{p}\bigl{(}(1+p^{e}u)^{p-1}-1\bigr{)}=e$, while for $p=2$ we have $v_{2}(a-1)=e$. In the standard notation $r_{p}(a):=v_{p}(a^{p-1}-1)-1$ (for $v_{p}(a)=0$) this yields
$r_{p}(a)=e-1,\qquad r_{p}(a^{\prime})\geq e.$
In the explicit bases of Jakhar–Khanduja–Sangwan, all pp–denominator exponents $k_{p,m}$ and the reduced linear corrections $\beta_{m}\bmod p^{k_{p,m}}$ depend only on $d_{p}(a):=\min\{r_{p}(a),e\}$ (together with $n,m$). Thus $d_{p}(a)=e-1$ while $d_{p}(a^{\prime})=e$ forces a change in the pp–part of the shape (see [JKS21, Theorem 1.6 and Theorem 1.8]). This proves (ii).
By (i) the pp–component of the shape has period $p^{e_{p}+1}$ for each $p\mid n$. By the Chinese Remainder Theorem the global shape is periodic modulo the product $M(n)=\prod_{p\mid n}p^{e_{p}+1}=n\cdot\operatorname{rad}(n)$. Conversely, if $N$ is any period, then for each fixed $p\mid n$ the congruence $a\equiv a^{\prime}\pmod{N}$ must imply $\mathbf{S}_{p}(a)=\mathbf{S}_{p}(a^{\prime})$; by (ii) this forces $p^{e_{p}+1}\mid N$. Since this holds for all $p\mid n$, we get $M(n)\mid N$. This proves (iii). ∎
@@
Separating the $p\mid n$ denominator profile from the factors $C_{m}(a)$ supported on primes dividing $a$ is standard in modern explicit constructions and is exactly the viewpoint in [JKS21]. For the restriction to primes $p\mid n$ in denominators see Remark 2.4, and for $n^{2}$–periodicity in small degrees see [GR17]. For the family $x^{p^{r}}-a$ and variants (with or without $(a,n)=1$), Newton–polygon methods compatible with our definitions appear in [Kch24].
By Theorem 3.1, $\mathbf{S}_{p}(a)$ depends only on $a\bmod p^{e+1}$; there are only $p^{e+1}$ classes. For each class, apply the order–$e$ Newton–polygon routine to $x^{n}-a$ to read off $k_{p,m}$ and the reduced $\beta_{m}$ (unique modulo $p^{k_{p,m}}$) via the pp–integral–basis extraction. It follows that there are finitely many possible pp–local shapes, and the map
@@
factors through $a\bmod p^{e+1}$. In particular one can enumerate all pp–local shapes by running Ore–Montes up to order $e$ for the $p^{e+1}$ residue classes.
@@
Fix $n$. There exists a finite lookup table $\mathcal{T}(n)$, indexed by the residue classes $a\bmod M(n)$, such that for any $n$th–power–free $a$ the shape of an integral basis of $\mathcal{O}_{K_{a}}$ is read off directly from the entry $a\bmod M(n)]\in\mathcal{T}(n)$. The table can be filled purely locally, prime–by–prime at $p\mid n$, by running higher Newton polygons up to order $e_{p}$.
@@
By Theorem 3.1 the pp–part of the shape depends only on $a\bmod p^{e_{p}+1}$. Thus $\mathcal{T}(n)$ is the Cartesian product of the local tables at $p\mid n$ listing, for each residue class modulo $p^{e_{p}+1}$, the exponents $k_{p,m}$ and the reduced corrections $\beta_{m}\bmod p^{k_{p,m}}$; these are exactly what the Ore–Montes output provides. The global shape is obtained by Chinese remaindering of the local choices. ∎
@@
Under $\gcd(n,\Delta_{L})=1$ there is no wild interaction at $\mathfrak{p}\mid p$, and the Ore–Montes computation of the $\mathfrak{p}$–integral basis runs identically to the absolute one. At order $t\leq e_{p}$ the residual polynomials use coefficients known modulo $p^{t+1}$, hence the local type is determined by $a\bmod p^{e_{p}+1}$ (cf. the index theorem and the pp–integral–basis extraction stated in [GMN12]). Gluing over all $\mathfrak{p}\mid p$ and all $p\mid n$ gives the asserted periodicity modulo $M(n)$. ∎
@@ 4. Applications of the minimal modulus
4.1. Sharp local monogeneity tests at primes $p\mid n$
@@
Suppose $p\parallel n$ and $v_{p}(a)=0$. Then the pp–local obstruction to a power $\mathbb{Z}$–basis is governed by the single congruence
$a^{p-1}\equiv 1\pmod{p^{2}}$.
More precisely, fixing $a\bmod p$ with $a^{p-1}\equiv 1\pmod{p}$, the order–$1$ type is pp–regular (hence the local index contribution at $p$ vanishes) iff $a^{p-1}\not\equiv 1\pmod{p^{2}}$; both outcomes occur among lifts of the same residue modulo $p$.
@@
When $e_{p}=1$, the entire pp–part of the shape depends on $d_{p}=\min\{r_{p},1\}$ with $r_{p}=v_{p}(a^{p-1}-1)-1$. If $a^{p-1}\not\equiv 1\pmod{p^{2}}$ then $r_{p}=0$ and the order–$1$ polygon is pp–regular (no pp–denominators). If $a^{p-1}\equiv 1\pmod{p^{2}}$ then $r_{p}\geq 1$, the order–$1$ type changes and $k_{p,m}$ jump accordingly, as encoded by the Jakhar–Khanduja–Sangwan construction. The two cases occur among lifts by Hensel’s lemma (take lifts of a fixed primitive (p1)(p-1)st root of unity). The regularity and index–zero conclusion at order 1 is the standard Ore–Dedekind criterion. ∎
@@
so raising $t$ to $t+1$ subtracts precisely $2\,n_{p}\,p^{\,e-(t+1)}$. The existence of such $a,a^{\prime}$ with fixed residue modulo $p^{e}$ and $t_{p}$ differing by one is ensured by Proposition 4.3. (We are specializing [JKS21, Theorem. 1.A] to a fixed $p^{e}\,\|\,n$.)
@@ 4.2. Counting and density of shape classes
We begin by stating (and proving) a precise form that cleanly separates the dependence on the residue class.
@@
Theorem 4.6 (Equidistribution with explicit rr–dependence).
Fix $q\geq 1$, $n\geq 2$, and a residue r(modq)r\ ($\mathrm${mod}\ q). Let
$\mathcal{F}_{n}(X;q,r)\ :=\ \#\{\,|a|\leq X:\ a\equiv r\pmod{q},\ a\text{ is $n$th–power–free}\,\}.$
Then, as $X\to\infty$,
$\mathcal{F}_{n}(X;q,r)\ =\ \frac{2X}{q}\,\frac{1}{\zeta(n)}\,\prod_{p\mid q}\Bigl{(}1-\frac{1}{p^{\,n}}\Bigr{)}^{-1}\ \prod_{p^{\alpha}\parallel q}\Bigl{(}1-\mathbf{1}_{\{v_{p}(r)\geq\alpha\}}\,p^{\alpha-n}\Bigr{)}\ +\ O\!\bigl{(}X^{1/n}\bigr{)}.$
@@
  • If $r$ is not nn–admissible, then $\mathcal{F}_{n}(X;q,r)=O(1)$ (indeed eventually 0).
  • If $r$ is strictly nn–admissible then the main term simplifies to
    $\mathcal{F}_{n}(X;q,r)\ =\ \frac{2X}{q}\,\frac{1}{\zeta(n)}\,\prod_{p\mid q}\Bigl{(}1-\frac{1}{p^{\,n}}\Bigr{)}^{-1}\ +\ O\!\bigl{(}X^{1/n}\bigr{)},$
@@
Write the indicator of $n$th–power–freeness by Möbius inversion:
$1_{\text{$n$–free}}(a)=\sum_{d^{n}\mid a}\mu(d).$
Hence
$\mathcal{F}_{n}(X;q,r)=\sum_{d\geq 1}\mu(d)\cdot\#\bigl{\{}|a|\leq X:\ a\equiv r\ (\mathrm{mod}\ q),\ d^{n}\mid a\bigr{\}}.$
For fixed $d$, set $g=\gcd(d^{n},q)$ and $L=\mathrm{lcm}(d^{n},q)=qd^{n}/g$. The two congruences are simultaneously solvable iff $g\mid r$, in which case they define a single class modulo $L$. Thus the inner count equals $\tfrac{2X}{L}+O(1)$ if $g\mid r$, and 0 otherwise. We obtain
$\mathcal{F}_{n}(X;q,r)=\frac{2X}{q}\sum_{d\geq 1}\mu(d)\frac{g}{d^{n}}\,\mathbf{1}_{\{g\mid r\}}\ +\ O\!\Bigl{(}\sum_{d\leq X^{1/n}}1\Bigr{)},$
and and the error is $O\!\big{(}\sum_{d\leq X^{1/n}}1\big{)}=O(X^{1/n})$.
The sum is multiplicative in $d$. We used that only $\nu_{p}(d)\in\{0,1\}$ contribute (since $\mu(p^{k})=0$ for $k\geq 2$) and that $g=\gcd(d^{n},q)=\prod_{p^{\alpha}\,\|\,q}p^{\min\{n\nu_{p}(d),\,\alpha\}}$. Writing $q=\prod p^{\alpha_{p}}$ and $d=\prod p^{\nu_{p}(d)}$, one has

### v2

@@ 3. Main theorems
Fix a prime $p\mid n$ with $p^{e}\parallel n$ and write $v=v_{p}$. Recall the invariants
$r_{p}(a)=\begin{cases}v_{p}(a^{p-1}-1)-1,&v_{p}(a)=0,\\ -1,&v_{p}(a)>0,\end{cases}\qquad d_{p}(a)=\min\{r_{p}(a),e\},$
and the denominator exponents $k_{p,m}$ defined in §2. By [JKS21, Theorem 1.6 and Theorem 1.8] (under hypothesis (H)), the pp–denominator exponents $k_{p,m}$ and the reduced correction terms $\beta_{m}\bmod p^{k_{p,m}}$ in the basis of Theorem 2.1 are determined by $p,n,m$ and the single invariant $d_{p}(a)$ (together with the branch $v_{p}(a)=0$ or $v_{p}(a)>0$).
(i) Local determinacy. Assume $a\equiv a^{\prime}\pmod{p^{e+1}}$. Then $v_{p}(a)>0$ if and only if $v_{p}(a^{\prime})>0$. If $v_{p}(a)>0$, then $r_{p}(a)=r_{p}(a^{\prime})=-1$ and hence $d_{p}(a)=d_{p}(a^{\prime})=-1$; therefore $k_{p,m}=0$ for all $m$. By Remark 2.3 the reduced correction data are trivial, so $\mathbf{S}_{p}(a)=\mathbf{S}_{p}(a^{\prime})$.
Assume now $v_{p}(a)=0$ (hence also $v_{p}(a^{\prime})=0$). From $a\equiv a^{\prime}\pmod{p^{e+1}}$ we have $a^{p-1}\equiv(a^{\prime})^{p-1}\pmod{p^{e+1}}$. Therefore
$\min\bigl\{v_{p}(a^{p-1}-1),\,e+1\bigr\}=\min\bigl\{v_{p}((a^{\prime})^{p-1}-1),\,e+1\bigr\}.$
It follows that
$d_{p}(a)=\min\{\,v_{p}(a^{p-1}-1)-1,\,e\,\}=\min\{\,v_{p}((a^{\prime})^{p-1}-1)-1,\,e\,\}=d_{p}(a^{\prime}).$
Since the exponents $k_{p,m}$ are defined purely from $d_{p}(a)$ (together with $p,n,m$), we get $(k_{p,m})_{m=1}^{n-1}=(k^{\prime}_{p,m})_{m=1}^{n-1}$. Finally, in the construction underlying [JKS21, Theorem 1.6], each reduced class $\beta_{m}\bmod p^{\,k_{p,m}}$ is computed from congruence data modulo $p^{\,k_{p,m}+1}$ (in particular from an inverse of $a$ modulo $p^{\,k_{p,m}+1}$). Since $k_{p,m}\leq e$ and $a\equiv a^{\prime}\pmod{p^{e+1}}$, the required congruence data agree for $a$ and $a^{\prime}$, hence the reduced correction terms coincide. Therefore $\mathbf{S}_{p}(a)=\mathbf{S}_{p}(a^{\prime})$.
(ii) Local sharpness. Take $p$-adic units
$a=1+p^{e}u,\qquad a^{\prime}=1+p^{e+1}u^{\prime},\qquad p\nmid uu^{\prime}.$
By replacing $a$ and $a^{\prime}$ by congruent integers if necessary, we may (and do) assume that $a$ and $a^{\prime}$ are squarefree and coprime to $n$; in particular they are $n$th–power–free, satisfy hypothesis (H), and $x^{n}-a$, $x^{n}-a^{\prime}$ are irreducible over $\mathbb{Q}$. (For example, each primitive residue class modulo $p^{e+1}$ contains infinitely many squarefree integers.) Then $a\equiv a^{\prime}\pmod{p^{e}}$ but $a\not\equiv a^{\prime}\pmod{p^{e+1}}$. For odd $p$, the lifting-the-exponent lemma gives $v_{p}(a^{p-1}-1)=e$ and $v_{p}((a^{\prime})^{p-1}-1)\geq e+1$; for $p=2$ this is the same statement with $a^{p-1}-1=a-1$. Thus $r_{p}(a)=e-1$ while $r_{p}(a^{\prime})\geq e$, hence $d_{p}(a)=e-1$ and $d_{p}(a^{\prime})=e$. Let $m_{0}:=n-n/p^{e}$. By definition of $k_{p,m}$ we have $k_{p,m_{0}}=e-1$ but $k^{\prime}_{p,m_{0}}=e$, so $\mathbf{S}_{p}(a)\neq\mathbf{S}_{p}(a^{\prime})$. This proves (ii).
(iii) Global minimal period. Part (i) shows that for each $p\mid n$ the local shape depends only on $a\bmod p^{e_{p}+1}$. By the Chinese Remainder Theorem, the global shape depends only on $a\bmod\prod_{p^{e_{p}}\parallel n}p^{e_{p}+1}=M(n)$, so $M(n)$ is a period. Conversely, if $N$ is any modulus such that $a\equiv a^{\prime}\pmod{N}$ implies $\mathbf{S}(a)=\mathbf{S}(a^{\prime})$ for all $n$th–power–free $a,a^{\prime}$, then for each fixed $p\mid n$ it implies $\mathbf{S}_{p}(a)=\mathbf{S}_{p}(a^{\prime})$. By (ii) this forces $p^{e_{p}+1}\mid N$ for all $p\mid n$, hence $M(n)\mid N$. ∎
@@
Separating the $p\mid n$ denominator profile from the factors $C_{m}(a)$ supported on primes dividing $a$ is standard in modern explicit constructions and is exactly the viewpoint in [JKS21]. For the restriction to primes $p\mid n$ in denominators see Remark 2.4, and for $n^{2}$–periodicity in small degrees see [GR17]. For the family $x^{p^{r}}-a$ and variants (with or without $(a,n)=1$), see [KCH24].
By Theorem 3.1, $\mathbf{S}_{p}(a)$ depends only on $a\bmod p^{e+1}$; there are only $p^{e+1}$ residue classes. For each class, the corresponding exponents $k_{p,m}$ and reduced corrections $\beta_{m}\bmod p^{k_{p,m}}$ can be read off from the explicit construction of Jakhar–Khanduja–Sangwan [JKS21, Theorem 1.6 and Theorem 1.8]. It follows that there are only finitely many possible pp–local shapes, and the map
@@
factors through $a\bmod p^{e+1}$.
@@
Fix $n$. There exists a finite lookup table $\mathcal{T}(n)$, indexed by the residue classes $a\bmod M(n)$, such that for any $n$th–power–free $a$ the shape of an integral basis of $\mathcal{O}_{K_{a}}$ is read off directly from the entry $a\bmod M(n)]\in\mathcal{T}(n)$. The table can be filled purely locally, prime–by–prime at $p\mid n$, using the explicit local recipes in [JKS21] for residue classes modulo $p^{e_{p}+1}$.
@@
By Theorem 3.1 the pp–part of the shape depends only on $a\bmod p^{e_{p}+1}$. Thus $\mathcal{T}(n)$ is the Cartesian product of the local tables at $p\mid n$ listing, for each residue class modulo $p^{e_{p}+1}$, the exponents $k_{p,m}$ and the reduced corrections $\beta_{m}\bmod p^{k_{p,m}}$; these are exactly what the explicit construction in [JKS21, Theorem 1.6 and Theorem 1.8] provides. The global shape is obtained by Chinese remaindering of the local choices. ∎
@@
Since $\gcd(n,\Delta_{L})=1$, every rational prime $p\mid n$ is unramified in $L$, so the local behavior at each prime $\mathfrak{p}\mid p$ is tame over $L$. The local denominator exponents and reduced correction terms in an $\mathcal{O}_{L}$–integral basis of $L(\sqrt[n]{a})/L$ at $\mathfrak{p}\mid p$ are governed by the same invariant $d_{p}(a)$ as in the absolute case (cf. [JKS21]), hence are determined by $a\bmod p^{e_{p}+1}$. Gluing over all $\mathfrak{p}\mid p$ and all $p\mid n$ yields periodicity modulo $M(n)$. ∎
@@ 4. Applications of the minimal modulus
4.1. Sharp local index tests at primes $p\mid n$
@@
Suppose $p\parallel n$ and $v_{p}(a)=0$ (equivalently, $(a,p)=1$). Then
$p\mid[\mathcal{O}_{K_{a}}:\mathbb{Z}[\theta]]\quad\Longleftrightarrow\quad a^{p-1}\equiv 1\pmod{p^{2}}$.
Equivalently, $\mathbb{Z}[\theta]$ is pp–maximal if and only if $a^{p-1}\not\equiv 1\pmod{p^{2}}$.
@@
When $p\parallel n$ (so $e_{p}=1$) and $v_{p}(a)=0$, we have
$d_{p}(a)=\min\{\,v_{p}(a^{p-1}-1)-1,\,1\,\}\in\{0,1\}.$
If $a^{p-1}\not\equiv 1\pmod{p^{2}}$ then $v_{p}(a^{p-1}-1)=1$, so $d_{p}(a)=0$ and the explicit basis in Theorem 2.1 has no pp–power denominators; equivalently, $v_{p}([\mathcal{O}_{K_{a}}:\mathbb{Z}[\theta]])=0$. If $a^{p-1}\equiv 1\pmod{p^{2}}$ then $v_{p}(a^{p-1}-1)\geq 2$, so $d_{p}(a)=1$ and the explicit basis has a nontrivial pp–denominator, hence $v_{p}([\mathcal{O}_{K_{a}}:\mathbb{Z}[\theta]])>0$. ∎
In this paper we work in the fixed coordinates $\{1,\theta,\dots,\theta^{n-1}\}$ and in the normalized form where all factors supported on primes dividing $a$ are absorbed into the numerators (via the $C_{m}(a)$). In this normalization, the only primes that can occur in the denominators are those dividing $n$ (see Remark 2.4). Accordingly, Proposition 4.1 concerns the pp–part of the index $\mathcal{O}_{K_{a}}:\mathbb{Z}[\theta]]$ at primes $p\mid n$; it is not meant to be a criterion involving primes dividing $a$ (indeed we assume $v_{p}(a)=0$ here).
Remark 4.2 (Index vs. monogeneity).
Proposition 4.1 is a criterion for the pp–adic contribution to the index $\mathcal{O}_{K_{a}}:\mathbb{Z}[\theta]]$, i.e. for pp–maximality of the power order $\mathbb{Z}[\theta]$. It does not decide whether $K_{a}$ is monogenic (whether there exists some $\alpha\in\mathcal{O}_{K_{a}}$ with $\mathcal{O}_{K_{a}}=\mathbb{Z}[\alpha]$).
@@
so raising $t$ to $t+1$ subtracts precisely $2\,n_{p}\,p^{\,e-(t+1)}$. The existence of such $a,a^{\prime}$ with fixed residue modulo $p^{e}$ and $t_{p}$ differing by one is ensured by Proposition $4.$ (We are specializing [JKS21, Theorem. 1.A] to a fixed $p^{e}\,\|\,n$.)
@@ 4.2. Counting and density of shape classes
By Theorem 3.1, for fixed $n$ the (global) shape $\mathbf{S}(a)$ depends only on the residue class of $a$ modulo $M(n)$. Therefore, any question of the form “how often does a given shape (or a union of shapes) occur as $|a|\leq X$?” reduces to counting those integers $a$ with $|a|\leq X$ that lie in prescribed congruence classes modulo $M(n)$ and are $n$th-power-free (because our fields $K_{a}=\mathbb{Q}(\sqrt[n]{a})$ are only defined for $n$th-power-free parameters). This is the analytic input behind our later shape density statements. Recall that an integer $a\neq 0$ is $n$th-power-free if no prime $p$ satisfies $p^{n}\mid a$. Equivalently, in the prime factorization $|a|=\prod_{p}p^{\alpha_{p}}$ one has $\alpha_{p}<n$ for every $p$. The sign of $a$ plays no role for $n$th-power-freeness.
We use a standard counting function for $n$th–power–free integers in arithmetic progressions. Fix integers $q\geq 1$ and r(modq)r\ ($\mathrm${mod}\ q). We define
$\mathcal{F}_{n}(X;q,r):=\#\bigl\{\,a\in\mathbb{Z}:\ 0<|a|\leq X,\ a\equiv r\ (\mathrm{mod}\ q),\ a\ \text{is $n$th-power-free}\,\bigr\}.$
Thus $\mathcal{F}_{n}(X;q,r)$ counts $n$th–power–free integers in the arithmetic progression $a\equiv r\pmod{q}$, up to height $X$ (counting both positive and negative integers; this explains the factor $2X$ that will appear in main terms). If $\mathcal{F}_{n}(X;q,r)$ grows linearly with $X$, then the leading constant is exactly the natural density (with respect to height $|a|$) of $n$th–power–free integers in that progression. Concretely, whenever the limit exists,
$\delta_{n}(q,r)\ :=\ \lim_{X\to\infty}\frac{\mathcal{F}_{n}(X;q,r)}{2X}$
is the proportion of integers (ordered by $|a|$) that lie in the class $r\bmod q$ and are $n$th–power–free. Theorem 4.8 computes this density and makes explicit how it depends on $r$.
Not every residue class modulo $q$ can contain infinitely many $n$th–power–free integers. For example, if $q$ contains a prime power $p^{\alpha}$ with $\alpha\geq n$ and $r\equiv 0\pmod{p^{\alpha}}$, then every $a\equiv r\pmod{q}$ is divisible by $p^{n}$, hence cannot be $n$th–power–free (except possibly $a=0$, which is excluded by $n$th–power–freeness). The following definitions isolate precisely this obstruction
@@
Remark 4.7 (Terminology).
Definition 4.6 concerns $n$th–power–free integers in a general arithmetic progression modulo $q$. Later, for the specific modulus $q=M(n)$ coming from the local theory, we will often restrict to the smaller set of residue classes with $v_{p}(r)\leq 1$ for every $p\mid n$; to avoid confusion we will call these shape–admissible classes (Definition 4.11 below).
For a nonzero integer $a$, the indicator of $n$th–power–freeness is
$1_{\text{$n$--free}}(a)\ :=\begin{cases}1,&\text{if $p^{n}\nmid a$ for every prime $p$,}\\ 0,&\text{otherwise.}\end{cases}$
A classical identity (see, e.g., [TEN15, Ch. III]) is
(2) $1_{\text{$n$--free}}(a)\ =\ \sum_{d^{n}\mid a}\mu(d).$
Here $\mu$ is the Möbius function. One quick justification is to factor the right-hand side as an Euler product: since $\mu(d)=0$ unless $d$ is squarefree, we have
$\sum_{d^{n}\mid a}\mu(d)=\prod_{p}\Bigl(1+\mu(p)\,\mathbf{1}_{\{p^{n}\mid a\}}\Bigr)=\prod_{p}\Bigl(1-\mathbf{1}_{\{p^{n}\mid a\}}\Bigr),$
which equals $1$ if no prime power $p^{n}$ divides $a$, and equals 0 otherwise, exactly as desired. Our counting function $\mathcal{F}_{n}(X;q,r)$ ignores $a=0$ anyway, since 0 is not $n$th–power–free.
We now evaluate $\mathcal{F}_{n}(X;q,r)$ asymptotically, keeping track of the precise dependence on $r$.
Theorem 4.8 (Equidistribution with explicit $r$-dependence).
Fix $q\geq 1$, $n\geq 2$, and a residue class r(modq)r\ ($\mathrm${mod}\ q). Then, as $X\to\infty$,
(3) $\mathcal{F}_{n}(X;q,r)=\frac{2X}{q}\cdot\frac{1}{\zeta(n)}\,\prod_{p\mid q}\Bigl(1-\frac{1}{p^{\,n}}\Bigr)^{-1}\,\prod_{p^{\alpha}\parallel q}\Bigl(1-\mathbf{1}_{\{v_{p}(r)\geq\min(\alpha,n)\}}\,p^{\min(\alpha,n)-n}\Bigr)\;+\;O\!\bigl(X^{1/n}\bigr).$
@@
  • If $r$ is not $n$-admissible, then $\mathcal{F}_{n}(X;q,r)=O(1)$.
  • If $r$ is strictly $n$-admissible, then the $r$-dependent product in (3) equals $1$, and the main term simplifies to
    $\mathcal{F}_{n}(X;q,r)=\frac{2X}{q}\cdot\frac{1}{\zeta(n)}\,\prod_{p\mid q}\Bigl(1-\frac{1}{p^{\,n}}\Bigr)^{-1}\;+\;O\!\bigl(X^{1/n}\bigr),$
@@
By definition,
$\mathcal{F}_{n}(X;q,r)=\sum_{\begin{subarray}{c}0<|a|\leq X\\ a\equiv r\ (\mathrm{mod}\ q)\end{subarray}}1_{\text{$n$-free}}(a).$
It follows that
$\mathcal{F}_{n}(X;q,r)=\sum_{\begin{subarray}{c}0<|a|\leq X\\ a\equiv r\ (\mathrm{mod}\ q)\end{subarray}}\ \sum_{d^{n}\mid a}\mu(d).$
The inner condition $d^{n}\mid a$ forces $d\leq X^{1/n}$ (since $0<|a|\leq X$), so the double sum is finite; there is no convergence issue, and we may swap the order of summation:
$\mathcal{F}_{n}(X;q,r)=\sum_{1\leq d\leq X^{1/n}}\mu(d)\cdot\#\bigl\{\,0<|a|\leq X:\ a\equiv r\ (\mathrm{mod}\ q),\ d^{n}\mid a\,\bigr\}.$
Fix $d\leq X^{1/n}$. We must count $a$ with
$a\equiv r\ (\mathrm{mod}\ q)\qquad\text{and}\qquad a\equiv 0\ (\mathrm{mod}\ d^{n}).$
Let
$g:=\gcd(q,d^{n}),\qquad L:=\mathrm{lcm}(q,d^{n})=\frac{q\,d^{n}}{g}.$
A standard compatibility criterion says that the two congruences are simultaneously solvable if and only if $r\equiv 0\ (\mathrm{mod}\ g)$, i.e. $g\mid r$. Indeed, any solution must be congruent to 0 modulo $g$, hence $r\equiv a\equiv 0\ (\mathrm{mod}\ g)$. Conversely, if $g\mid r$, then dividing the second congruence by $g$ reduces it to a single class modulo $q/g$, and since $\gcd(q/g,\,d^{n}/g)=1$, the Chinese Remainder Theorem produces a unique solution modulo $L$. Hence there is either:
  • no solution if $g\nmid r$, or
  • exactly one residue class $a\equiv a_{0}\ (\mathrm{mod}\ L)$ if $g\mid r$.
In the solvable case, the number of integers $a$ with $0<|a|\leq X$ in one fixed class modulo $L$ is
$\frac{2X}{L}+O(1),$
since such integers form a two-sided arithmetic progression with step $L$. Therefore
#{ $\bigl\{\,0<|a|\leq X:\ a\equiv r\ (\mathrm{mod}\ q),\ d^{n}\mid a\,\bigr\}=\mathbf{1}_{\{g\mid r\}}\Bigl(\frac{2X}{L}+O(1)\Bigr)=\mathbf{1}_{\{g\mid r\}}\Bigl(\frac{2X}{q}\cdot\frac{g}{d^{n}}+O(1)\Bigr).$
Plugging this into the $d$-sum yields
$\mathcal{F}_{n}(X;q,r)=\frac{2X}{q}\sum_{1\leq d\leq X^{1/n}}\mu(d)\,\frac{g}{d^{n}}\,\mathbf{1}_{\{g\mid r\}}\;+\;O\!\Bigl(\sum_{d\leq X^{1/n}}1\Bigr).$
Since $\sum_{d\leq X^{1/n}}1\ll X^{1/n}$, the error term is $O(X^{1/n})$.
Now we analyze
$S(q,r):=\sum_{d\geq 1}\mu(d)\,\frac{\gcd(d^{n},q)}{d^{n}}\,\mathbf{1}_{\{\gcd(d^{n},q)\mid r\}}.$
Only squarefree $d$ contribute because $\mu(d)=0$ otherwise, so write $d=\prod_{p}p^{\nu_{p}(d)}$ with $\nu_{p}(d)\in\{0,1\}$. Also write $q=\prod_{p}p^{\alpha_{p}}$ where $\alpha_{p}=v_{p}(q)$. Then
$\gcd(d^{n},q)=\prod_{p}p^{\min(n\nu_{p}(d),\alpha_{p})}.$
Consequently, the condition $\gcd(d^{n},q)\mid r$ is equivalent to requiring, for every prime $p$ with $\nu_{p}(d)=1$, that
$v_{p}(r)\geq\min(n,\alpha_{p}).$
Moreover,
$\frac{\gcd(d^{n},q)}{d^{n}}=\prod_{p:\,\nu_{p}(d)=1}p^{\min(n,\alpha_{p})-n}.$
This shows that the sum $S(q,r)$ factors as a product of independent local contributions over primes. For a given prime $p$, only $\nu=0$ and $\nu=1$ occur:
$\sum_{\nu\in\{0,1\}}\mu(p^{\nu})\,p^{\min(n\nu,\alpha_{p})-n\nu}\,\mathbf{1}_{\{\min(n\nu,\alpha_{p})\leq v_{p}(r)\}}=1-\mathbf{1}_{\{v_{p}(r)\geq\min(n,\alpha_{p})\}}\,p^{\min(n,\alpha_{p})-n}.$
If $p\nmid q$ then $\alpha_{p}=0$ and the local factor becomes $1-p^{-n}$.
Therefore,